### v1

# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

Abstract
We propose to solve the SUSY $\mu$-problem in the framework of modular flavor symmetry. The bare $\mu$-term is forbidden by modular symmetry and the effective $\mu$-term is re-generated after the modulus field is stabilized. We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield. The presence of $X_{nl}$ contributions can possibly slightly change the original stabilized modulus VEV to a value located in the vicinity of some fixed points. Natural value of $\mu_{eff}$, which is much smaller than the SUSY scale, can either be the consequence of expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), or be the joint consequence of suppression from both powers of $q$ [or $(2\Im\tau)^{-1}$] and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
1 Introduction
Low energy supersymmetry (SUSY) is well motivated theoretically and regarded as one of the most appealing extensions of the Standard Model. Not only the quadratic divergences related to the Higgs mass can be eliminated, but also successful gauge coupling unification can be realized in the framework of low energy SUSY. Especially, the observed 125 GeV scalar, discovered by both the ATLAS ATLAS and CMS CMS collaborations of LHC, falls miraculously within the narrow $115\sim 135$ GeV ’window’ required by minimal supersymmetric standard model (MSSM) and it can be seen as a triumph of low energy SUSY.
Although the framework of low energy SUSY is attracting, it still has several unsatisfactory features. In addition to the problem of null observation of conclusive signals for superparticles at the LHC, the MSSM is also bothered by the SUSY $\mu$-problem: the $\mu$-parameter is a SUSY-respecting parameter appearing in the superpotential, while it is required to be roughly of order the soft SUSY breaking scale to allow a Higgs VEV of order 174 GeV without too much miraculous cancelation.
There are many proposals to tackle the $\mu$-problem, for example, by singlet extension to next-to-minimal supersymmetric standard model (NMSSM) or by non-renormalizable Lagrangian terms in the superpotential via Kim-Nilles mechanism KN (or in the Kahler potential via Giudice-Masiero mechanism GM ). Most solutions consist of two steps Bae:2019dgg : First, the appearance of $\mu$-parameter is forbidden, usually by some discrete or continuous symmetry; Second, the $\mu$ term is re-generated via some mechanism such as symmetry breaking. One can see in Bae:2019dgg some detailed realizations of the two steps for the $\mu$-problem.
We would like to tackle the $\mu$-problem with a similar but different approach. The symmetry adopted to forbid the $\mu$-parameter is a modular flavor symmetry. Besides, the re-generation of $\mu$ term is realized after modular symmetry breaking with the modulus stabilization mechanism accompanied by typical SUSY breaking contributions. Naturally suppressed $\mu$-parameter can thus be realized.
Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis. Modular symmetry can naturally arise in string theory, for example, (exact) moduli spaces of Calabi-Yau threefolds, like the quintic and fairly generically in large classes of elliptic Calabi-Yau threefolds. Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the standard model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings. For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt . The modular symmetry can constrain stringently the superpotential as the relevant Yukawa couplings are modular forms. The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$). In modular flavor symmetry framework, the masses and mixings for the quarks and leptons can be predicted with very few input parameters. Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been adopted in various studies to explain the low energy flavor structures of the Standard Model. Besides, modular flavor symmetries can combine with the GUT framework to increase greatly the prediction ability of the models, such as the combination of $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking, it can also be the only source of CP symmetry breaking. The VEV of the modulus field $\tau$ will in general break the modular symmetry completely except three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain. The modular forms, thus the Yukawa coupling, and the fermion mass matrices will be fixed after the modular flavour symmetry is fully or partially broken by VEV of $\tau$. From a top-down perspective, the value of modulus VEV can be determined dynamically with modular-invariant scalar potential in 4d $N=1$ supergravity, which is also crucial to give nonzero masses for modulus field to render it phenomenological acceptable.
Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy. In the presence of additional ”matter”-like fields, it is possible to obtain dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induces a SUSY breaking vacuum shift to a tiny positive vacuum energy. The presence of the anti-brane like sources can be described in terms of a nilpotent goldstino superfield KS in a supersymmetric effective field. Although SUSY is non-linearly realized, the couplings of chiral, gauge and gravity superfields coupled to the constrained nilpotent goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action.
This paper is organized as follows. In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry. In Sec 3, we discuss the solution of SUSY $\mu$-problem with modular symmetry. In Sec 4, we discuss the stabilization of singlet modulus field, with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield. In Sec 5, we discuss the re-generation of naturally small $\mu_{eff}$ after modulus stabilization. Sec 6 contains our conclusions.
2 Modular $\Gamma_{N}$ symmetry
The full modular group is the group of linear fraction transformations acting on the complex modulus $\tau$ in the upper half complex plane
$\gamma=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right):\tau\longmapsto\gamma\tau=\frac{a\tau+b}{c\tau+d},~{}~{}\text{with}~{}~{}a,b,c,d\in\mathbb{Z},~{}~{}ad-bc=1,~{}~{}~{}{\rm Im}\tau>0\,.$ (1)
It can be generated by two generators $S$ and $T$
$S:\tau\mapsto-\frac{1}{\tau},~{}~{}~{}~{}\quad T:\tau\mapsto\tau+1\,,$ (2)
which satisfy
$S^{4}=1,~{}~{}(ST)^{3}=1,~{}~{}S^{2}T=TS^{2}.$ (3)
The two generators can be represented by the following $SL(2,Z)$ matrices,
$S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~{}~{}~{}\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)\,.$ (4)
Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined by $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, with $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ defined by
$\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}N\geq 3\\ \Gamma(N)/\{\pm 1\},~{}~{}~{}N=2\end{array}\right.$ (7)
with $\Gamma(N)$ the principal congruence subgroup of level $N$
$\Gamma(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL(2,\mathbb{Z}),~{}b=c=0\,(\mathrm{mod~{}N}),a=d=1\,(\mathrm{mod~{}N})\right\}\,.$ (8)
Finite modular groups $\Gamma_{N}$ can be obtained by further imposing the condition $T^{N}=1$ in addition to $S^{2}=(ST)^{3}=1$.
Modular form $f(\tau)$ of weight $2k$ and level $N$ is a holomorphic function of the complex modulus $\tau$ with well-defined transformation properties under the action of ${\Gamma}(N)$
$f\left(\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{2k}f(\tau)~{},~{}~{}~{}\forall~{}\gamma\equiv\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in{\Gamma}(N)\,.$ (9)
Modular forms of weight $2k$ and level $N$ span a linear space of finite dimension. They can be arranged into some modular multiplets $Y_{\mathbf{r}}^{2k}$ that transform as irreducible representation $\mathbf{r}$ of the finite modular group $\Gamma_{N}$ up to automorphy factor Feruglio:2017spp , i.e.
$Y_{\mathbf{r}}^{(2k)}(\gamma\tau)=(c\tau+d)^{2k}\rho_{\mathbf{r}}(\gamma)Y_{\mathbf{r}}^{(2k)},~{}~{}~{}~{}\forall~{}\gamma\in\overline{\Gamma}\,,$ (10)
where $\gamma$ is the representative element of the coset $\gamma\overline{\Gamma}(N)$ in $\Gamma_{N}$, and $\rho_{\mathbf{r}}(\gamma)$ is the representation matrix of the element $\gamma$ in the irreducible representation $\mathbf{r}$. The actions of modular group on supermultiplets read
$\phi_{i}(\gamma\tau)=(c\tau+d)^{-k_{\phi}}\rho^{\mathbf{r^{\prime}}}_{ij}(\gamma)\phi_{j}~{},~{}~{}~{}\forall\gamma\in\bar{\Gamma}~{},$ (11)
where $k_{\phi}$ denotes the modular weight of $\phi$, and $\rho$ is a unitary representation of the quotient group $\bar{\Gamma}_{N}$.
3 Natural $\mu$-term in modular flavor model
As noted in Bae:2019dgg , to solve the SUSY $\mu$ problem, the appearance of $\mu$-parameter in the superpotential should be forbidden by some symmetry. We propose that the superfields $H_{u}$, $H_{d}$ transform non-trivially under modular transformation so as that ordinary $\mu H_{u}H_{d}$ term is forbidden by modular symmetry.
The relevant Lagrangian can be given as
${\cal L}$ $\supseteq$ $\int d^{4}\theta\left[-3\log(2\Im\tau)+\frac{H_{u}^{\dagger}H_{u}}{(2\Im\tau)^{k_{u}}}+\frac{H_{d}^{\dagger}H_{d}}{(2\Im\tau)^{k_{d}}}+\cdots\right]$ (12)
$+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
with the coefficient for bilinear $H_{u}H_{d}$ term given by
$\mu=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ with $62k_{3}\in 24\mathbb{Z}+$ 6 to guarantee the modular invariance of
$e^{G}\equiv e^{K}|W|^{2}~{}.$ (14)
Here, $\alpha\sim{\cal O}(1)$ is a dimensionless parameter and $\Lambda_{SUSY}$ is typically of order the SUSY scale. The Dedekind $\eta$-function is a modular form with modular weight $1/2$ under modular transformation up to a phase, whose transformation under $\gamma$ can be given as
$\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~{},~{}~{}\forall\gamma\in\overline{\Gamma}~{}.$ (15)
with $\epsilon_{\eta}(\gamma)$ depending on $\gamma$ (but not on $\tau$) and being given in terms of the Rademacher $\Phi(\gamma)$ function
$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right].$ (16)
that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx .
The expansion of $\eta(\tau)$ function is
$\eta(\tau)$ $=$ $q^{1/24}\sum\limits_{n=-\infty}^{\infty}(-1)^{n}q^{n(3n-1)/2},$ (17)
$=$ $q^{1/24}\left[1+\sum\limits_{n=1}^{\infty}(-1)^{n}\left(q^{n(3n-1)/2}+q^{n(3n+1)/2}\right)\right],$
$=$ $q^{1/24}\left[1-q-q^{2}+q^{5}+q^{7}-q^{12}-\cdots\right],~{}~{}~{}~{}~{}q\equiv e^{i2\pi\tau}~{}.$
It is obvious that the Kahler potential for superfields with non-trivial modular weights, for example, the $H_{u}$ and $H_{d}$ superfields, are non-minimal. Therefore, after the moduli fields $\tau$ acquire a VEV by modulus stabilization mechanism, the $H_{u},H_{d}$ fields need to be normalized properly to get their canonical forms, which amounts to the replacements
$H_{u}\rightarrow H_{u}(2\Im\tau)^{k_{u}/2}~{},~{}~{}~{}H_{d}\rightarrow H_{d}(2\Im\tau)^{k_{d}/2}~{}.$ (18)
So, when the modulus field is fixed by modulus stabilization, the effective $\mu$-parameter will be re-generated to be
$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{u}+k_{d})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
We need to explain why such $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and be of order the soft SUSY breaking scale $m_{soft}$. We indicate that the smallness of $\mu_{eff}$ is the consequence of large suppression factor, either origin from some power of deviation of modulus VEV from typical fixed point or from the non-trivial power (that determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of some factor that related to the modulus VEV. All of the suppression factors are determined by the modular stabilization mechanism, with the stabilized modulus VEVs affected by SUSY breaking effects.
4 Stabilization of modulus field with SUSY breaking effects
To stabilize the modulus field, effective supergravity scalar potentials from weight $-3$ modular form superpotential can be used to determine the minimum. The modulus VEV can always be restricted to the fundamental domain $D_{H}$
${D}_{F}\equiv\left\{\tau\in\mathcal{H}:-\frac{1}{2}\leq\Re{\tau}<\frac{1}{2},\,|\tau|>1\right\}\cup\left\{\tau\in\mathcal{H}:-\frac{1}{2}<\Re\tau\leq 0,\,|\tau|=1\right\}\,,$ (20)
of the modular group, as any choice of the modulus VEV in the upper-half plane can be uniquely related to a $\tau\in D_{F}$ in the fundamental domain via a modular transformation, which is physically equivalent to such a choice. Discussions on the modulus stabilization with single modulus and multiple moduli can be found in Kobayashi:2019xvz ; NPP:2201.02020 ; PLNSR:2304.14437 ; King:2023snq ; King:2024ssx ; Ding:2024neh ; Higaki:2024jdk ; Kobayashi:2023spx ; Funakoshi:2024yxg .
To guarantee the modular invariance and Kahler transformation invariance of
$e^{G}\equiv e^{K}|W|^{2}~{},$ (21)
the superpotential for modulus should take the following form
$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}}.$ (22)
for the Kahler potential of the form
$K_{mod}=-3\log(2\Im\tau)~{},$ (23)
in the unit $M_{Pl}=1$ and $c_{1}M_{Pl}$ denotes the typical scale for the superpotential for dimensionless number $c_{1}$.
The modular invariant holomorphic function $H(\tau)$ is given by
$H(\tau)_{m,n}$ $=$ $\left(j(\tau)-1728\right)^{m/2}j(\tau)^{n/3}{\cal P}\left(j(\tau)\right),$ (24)
$\equiv$ $\left(\frac{G_{6}(\tau)}{\eta(\tau)^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\eta(\tau)^{12}}\right)^{n}{\cal P}\left(j(\tau)\right)~{},$
where ${\cal P}$ denotes a polynomial with respect to $j(\tau)$. The Klein j-function is defined as a modular form of zero weight
$j(\tau)$$\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\eta(\tau)^{24}},$ (25)
$\approx$ $744+q^{-1}+196884q+21493760q^{2}+864299970q^{3}+{\cal O}(q^{4}),$
with $j(i\infty)$ diverges, $j(\omega)=0$ and $j(i)=1728$.
We should note that, in such a stabilization mechanism, fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although may not be the global minimum. With proper SUSY breaking contributions, possibly from the ”matter”-like sector, the stabilized modulus value, which located at/near the fixed points $z_{FP}=\omega,i$, can slightly deviate from the original global minimum and possibly uplift the vacuum to dS (de Sitter) one. In fact, the classical 10d supergravity Maldacena-Nunez no-go result MN suggests that any constructions that lead to dS vacuum must include non-classical ingredients as D-branes, O-planes, or quantum corrections. To describe the scenario in which SUSY is linearly realized in a gravitational sector and non-linearly realized in the uplifting SUSY breaking sector, (for example, from anti-D branes) the formalism of non-linearly realized SUSY as formulated by KS can be used for such a system. It was proved in Burgess:2021juk with the non-linear SUSY realization scheme that the supergravity form is stable even if the supermultiplets in SUSY breaking sector are split badly enough to allow some of its members to be integrated out while the others are not.
When SUSY is spontaneously broken in supergravity, the goldstino is eaten by the gravitino via the super-Higgs mechanism. When such a breaking happens at low energies compared with the Planck mass, the goldstino couplings can be described by introducing a (constrained) independent superfield in the supergravity effective action. The goldstino provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. The emergence of the constraints may also be viewed as an effective field theory limit in which certain mass parameters become very large and the corresponding modes decouple.
Coupling of non-linearly realized SUSY with nilpotent goldstino $X_{nl}^{2}=0$ can be described by
${\cal L}\supseteq\int d^{4}\theta X_{nl}^{*}X_{nl}+\left[\int d^{2}\theta f_{0}X_{nl}+h.c.\right]~{},$ (26)
with the chiral superfield
$X_{nl}=\frac{\psi_{X}\psi_{X}}{2F_{X}}+\sqrt{2}\theta\psi_{X}+\theta^{2}F_{X}~{},$ (27)
is the IR limit of certain microscopic superfield. As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $X$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent goldstino. From the perspective of low energy effective theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for modulus stablization and vacuum uplifting discussions.
When coupled to supergravity, the most general form is given by
[…]
with
$K={X}^{*}_{nl}X_{nl}~{},~{}~{}~{}W=f_{0}X+W_{0}~{}.$ (29)
and $f_{0}=-F_{X}$.
To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
$K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{{X}^{*}_{nl}X_{nl}}{(2\Im\tau)^{k_{X}}}~{},$
$W$ $\supseteq$ $\left(f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
The scalar potential for $X$ and $\tau$ can be obtained from
$V=e^{K}\left[K^{\alpha\bar{\beta}}\nabla_{\alpha}W\nabla_{\bar{\beta}}W-3W\overline{W}\right],$ (31)
with
$\nabla_{\alpha}W=W_{\alpha}+K_{\alpha}W~{}.$ (32)
From the Kahler potential, we can calculate the inverse Kahler metric
$K^{\bar{\alpha}\beta}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{|X|^{2}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}X}{3}\\ -\frac{i(2\Im\tau)k_{X}{X}^{*}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}|X|^{2}}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
Here the indices $\alpha\sim(\tau,X_{nl})$ and the nilpotency of $X_{nl}$ is used in the calculation.
The scalar potential can be calculated to be
$V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (36)
$=$ $\frac{e^{\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left(1-\frac{k_{X}}{3}\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}\right)\left|\tilde{W}_{M}\tilde{c}_{M}+H(\tau)\tilde{W}_{T}\right|^{2}\right.$
$+$ $(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}|X_{nl}|^{2}}{3(2\Im\tau)^{k_{X}-1}}\right]\left|H(\tau)\tilde{W}_{X}\right|^{2}$
$+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}(\tilde{W}_{X}H(\tau))^{*}\left(\tilde{W}_{M}\tilde{c}_{M}+H(\tau)\tilde{W}_{T}\right)\right]$
-- $\left.3\left|\tilde{c}_{W}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
with
$\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
$\tilde{W}_{T}$$\equiv$ $k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\left[\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}+\frac{ic_{1}^{3}X_{nl}^{*}}{f_{0}(2\Im\tau)^{k_{X}+1}\eta(\tau)^{2k_{X}}}\right]~{},$
$\tilde{c}_{M}$ $=$ $f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}~{},$
$\tilde{W}_{X}$ $=$ $f_{0}\eta(\tau)^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{2(\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (37)
and the constraint $X_{nl}^{2}=0$ is used to eliminate some of the high order $X_{nl}$ terms.
We also use the expression for the derivative of Dedekind $\eta$ function, which is given by
$\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}~{},$ (38)
where the holomorphic Eisenstein series $G_{2}$, which is a quasi-modular form
$G_{2}(\gamma\tau)=(c\tau+d)^{2}G_{2}(\tau)-2\pi ic(c\tau+d)~{},$ (39)
is related to its weight two non-holomorphic counterpart $\hat{G}_{2}(\tau,\bar{\tau})$ as
$\hat{G}_{2}(\tau,{\tau}^{*})=G_{2}(\tau)-\frac{\pi}{\Im\tau}~{}.$ (40)
The generic complicate form of the scalar potential may obscure the discussions of the stabilization mechanism. So, we would like to concentrate on its typical simplification forms
  • Scenario with $k_{X}=0$:
    The scalar potential can be simplified to be
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (41)
    $=$ $\frac{e^{|X_{nl}|^{2}}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\tilde{c}_{M}\right|^{2}\right.$
    $+$ $\left.\left|H(\tau)\right|^{2}\left|f_{0}\left(1+{|X_{nl}|^{2}}\right)+{{X}^{*}_{nl}c_{1}^{3}}\right|^{2}-3\left|\tilde{c}_{M}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right.$
    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}{|X_{nl}|^{2}}+2\Re\left[f_{0}c_{1}^{3}{X}_{nl}^{*}\right]+|X_{nl}|^{2}c_{1}^{6}\frac{}{}\right)~{}$
    -- $\left.3\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\right\}~{}$
    $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}c_{1}^{3}\right|^{2}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{},$
    with
    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{},$ (42)
    Note the constraints $X_{nl}^{2}=\bar{X}_{nl}^{2}=0$ are imposed for the scalar potential.
    The modulus VEV can be determined by the extrema conditions of the scalar potential
    $\frac{\partial}{\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=\frac{\partial}{\partial\tau^{*}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=0~{},$
    and the minimum condition can be determined by positive definiteness of Hessian matrix in terms of
    $\frac{\partial^{2}}{\partial x^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}+\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
    $\frac{\partial^{2}}{\partial y^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}-\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
    $\frac{\partial^{2}}{\partial x\partial y}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $-2\Im\frac{\partial^{2}}{\partial\tau\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$ (43)
    with $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$, as well as those conditions for $X_{nl}$. It was argued in target:space ; Ding:2024neh that, modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensures that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
    $\left.\frac{\partial}{\partial x}V\right|_{x=\pm 1/2,0}=0~{},~{}~{}~{}\left.\frac{\partial}{\partial\rho_{\tau}}V\right|_{\rho_{\tau}=1}=0~{}.~{}~{}~{}$ (44)
    Then we will determine the extrema condition
    $\frac{\partial}{\partial\rho}V=\frac{\partial}{\partial\theta}V=0~{},$ (45)
    for $X_{nl}\equiv\rho e^{i\theta}$. It can be calculated that the extreme conditions for $X_{nl}$ reduce to the constraints
    $\frac{\partial}{\partial\theta}V=0~{},\Rightarrow\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{},$ (46)
    and
    $\frac{\partial}{\partial\rho}V=0~{},$$\Rightarrow$ $\left|H(\tau)\right|^{2}f_{0}^{2}=0~{},$ (49)
    $\Rightarrow$ $\left\{\begin{array}[]{l}\left|H(\tau)\right|^{2}=0,~{}~{}~{}~{}~{}~{}{\rm for}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}f_{0}\neq 0~{}.\\ ~{}~{}~{}~{}~{}~{}~{}f_{0}=0~{},~{}~{}~{}~{}{\rm and/or}~{}~{}~{}\left|H(\tau)\right|^{2}\neq 0.\end{array}\right.$
    So, we have the minimum conditions for $X_{nl}$
    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{},$ (50)
    in the case $f_{0}\neq 0$. Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes. So, the presence non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe.
    We will concentrate on the constraint $H(\tau)=0$ when $f_{0}\neq 0$. When $f_{0}=0$, the scalar potential for modulus fields will not be affected by $X_{nl}$ as the two sector decouples in the $k_{X}=0$ limit.
    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires either the condition $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0$ or ${\cal P}(j(\tau))=0$. Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively. As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, arbitrary complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all. So, we would like to focus on the constraint
    $\tilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0~{},$ (51)
    that satisfies $H(\tau)=0$. Such a constraint reduces to
    $H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))\right\}^{\prime}=0~{},$ (52)
    $\Rightarrow$ $j^{\prime}(\tau)\left[\frac{n}{3}j(\tau)^{n/3-1}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))+\frac{m}{2}j(\tau)^{n/3}[j(\tau)-1728]^{m/2-1}{\cal P}(j(\tau))\right.$
    $\left.+j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
    So $H^{\prime}(\tau)=0$ can trivially hold for $H(\tau)=0$.
    On the other hand, $H^{\prime}(\tau)=0$ can also be satisfied by the other two conditions
    $\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$ (53)
    and $j^{\prime}(\tau)=0$.
    We have the following discussions for the condition (53)
    • $n\neq 0$ and $m=0$ ($m\neq 0$ and $n=0$) or $mn\neq 0$:
      The condition (53) reduces to
      $\frac{d{\cal P}(j(\tau))}{{\cal P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)}~{},$ (54)
      which gives
      ${\cal P}(j(\tau))=C_{0}\left[j(\tau)\right]^{-n/3}~{}.$ (55)
      So, unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$. Similarly, we can deduce that $m\in 2\mathbb{Z}$ should be negative. Such a deduction can also hold for $m\neq 0$ with $n\neq 0$ case, which also requires $m=n=0$. As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
    • $m=0$ and $n=0$:
      With the constraints that $m=n=0$, the condition (53) reduces to
      $\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}=0~{},$ (56)
      which gives
      ${\cal P}(j(\tau))={\rm Constant}.$ (57)
      as a constant. As $H(\tau)=0$, the constant should be zero.
    So, the condition (53) requires $H(\tau)$ to be zero. As the constraint (53) is trivial, we would like to discuss the condition $j^{\prime}(\tau)=0$. From the derivative of the Klein j-function
    $\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~{},$ (58)
    we can see that its solutions are $\omega$ and $i$.
    So, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential can only be determined by $H(\tau)=0$ (and the extrema conditions from $X_{nl}$ are also satisfied by $z_{0}=\omega,i$). The corresponding vacuum is Minkowski type. As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are the local extrema. Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$. So, other minimum, which is determined by (44) to locate at the boundary of the fundamental domain or at the imaginary axis, is also possible.
    In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices. It can be seen that the energies are non-negative when $f_{0}\neq 0$.
    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$.
  • Scenario with small $X_{nl}$ (when $|X_{nl}|\rightarrow 0$):
    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the . constraint $X_{nl}^{2}=0$. The scalar potential can be simplified to
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (59)
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\left[\eta(\tau)\right]^{2k_{X}}\right]\right)\right.$
    $+$ $\left.\frac{(2\Im\tau)^{2}}{3}\left(2\Re\left[k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\tilde{W}_{E}\right]\right)H(\tau)c_{1}^{3}\tilde{W}_{M}^{*}\right]$
    $+$ $(2\Im\tau)^{k_{X}}\left|H(\tau)\right|^{2}\left(f_{0}^{2}\left|\eta(\tau)\right|^{4k_{X}}+2\Re\left[f_{0}\eta(\tau)^{2k_{X}}\frac{{X}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}\right]\right)$
    $+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}\left(f_{0}[\eta(\tau)]^{2k_{X}}H(\tau)\right)^{*}\tilde{W}_{M}c_{1}^{3}\right]$
    -- $\left.3\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\frac{}{}\right\}~{},$
    with
    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
    $\tilde{W}_{E}$$\equiv$ $\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}~{}.$ (60)
    It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$. So, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point can possibly be the minimum. The (otherwise stabilized) modulus VEV (without contributions from $X_{nl}$ sector) can be destabilized by non-vanishing $X_{nl}$. Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below. We will discuss the extrema conditions for the most general case in our subsequent studies.
5 Re-generation of naturally small $\mu_{eff}$ after modulus stabilization
It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw . One can easily seen from the modular invariance properties of scalar potential that the self-dual fixed points should be the local extrema of the scalar potential, since they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish. Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$ for the tree-level scalar potential, as the scalar potential in general diverges for $\tau\rightarrow i\infty$.
It was found in Novichkov:2021evw that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields). Similar ideas, augmented with additional suppression powers of $q$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis. We anticipate that, due to the SUSY breaking contributions from $X_{nl}$ sector, the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be pushed slightly away from the fixed points $\tau_{FP}=i,\omega\equiv e^{i2\pi/3}$. Such deviation can lead to the small parameters within the $\mu_{eff}$.
Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the scalar potential can still possibly lead to the stabilization of $\tau$ to a value with $\Im\tau_{0}$ being medium because of the contributions from $X_{nl}$. Then, for proper choice of $k_{3}$, the effective $\mu$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$ by a large suppression factor, which origins from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression factor from modular forms at $\tau_{0}$ (with medium $\Im\tau_{0}$).
5.1 Natural $\mu_{eff}$ with modulus Stabilization near fixed points $\omega\equiv e^{i2\pi/3}$
The Higgs superfields $H_{u},H_{d}$ in MSSM should always be singlets under modular flavor symmetry. They can be assigned with some singlet representations and non-trivial modular weights for their superpotential.
For modular $A_{4}$ symmetry, at the point $\omega$, the following modular forms
$Y_{\bf 1}^{(4)}(\omega)=Y_{\bf 1}^{(8)}(\omega)=Y_{\bf 1^{\prime}}^{(8)}(\omega)=Y_{\bf 1}^{(12)}(\omega)=Y_{\bf 1^{\prime}}^{(12)}(\omega)=Y_{\bf 1}^{(16)}(\omega)=0~{},$ (61)
vanish as well as $Y_{\bf 1^{\prime\prime}}^{(16)}(\omega)$ etc. The small deviation of modulus VEV away from $\omega$ by modulus stabilization mechanism can act as the desired suppression parameter for $\mu_{eff}$.
Expressions of modular forms with larger modular weights can be constructed from $Y_{\bf 3}^{(2)}\equiv(Y_{1},Y_{2},Y_{3})$ by
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~{},$
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})^{3}(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(16)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{4}~{}.$ (62)
From the expansion Petcov:2022fjf at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ satisfying $|\epsilon|\ll 0$) near $\omega$ point
$\frac{Y_{2}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $\omega\left(1+2.235i\epsilon-2.95717\epsilon^{2}+\cdots\right)~{},$
$\frac{Y_{3}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $-\frac{1}{2}\omega^{2}\left(1+4.47i\epsilon-10.9096\epsilon^{2}+\cdots\right),$ (63)
one can arrive at the expansion for $Y_{\bf 1}^{(16)}(z_{0})$ etc as
$Y_{\bf 1^{\prime}}^{(8)}(\omega+\epsilon)$ $\simeq$ $(10.5823+6.10969i)\epsilon-(39.9459-69.1884i)\epsilon^{2}+\cdots~{},$
$Y_{\bf 1}^{(16)}(\omega+\epsilon)$ $\simeq$ $1325.96\epsilon^{4}+18871.7i\epsilon^{5}+\cdots~{},$ (64)
with $Y_{1}(\omega)\approx 0.948674$.
[…]
$W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}~{},$ (65)
with $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6 and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$. When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{u}+k_{d})/2}~{},$ (66)
$=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
$\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~{}.$
So, to obtain much suppressed $\mu_{eff}$, we either need very large negative value $k_{3}$ or large $k_{r}$ to get large powers of $\epsilon$. However, it can be calculated that the suppression from $k_{3}$ power is not very efficient to get tiny number. So we choose to adopt large powers of $\epsilon$ for suppressed $\mu_{eff}$ with proper modular weights and representations for $H_{u},H_{d}$.
For example, we can adopt the choice
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime}},~{}6),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime\prime}},~{}7)~{},$ (67)
and the choice of modular forms $Y_{\bf 1}^{(16)}$ (or possibly $Y_{\bf 1}^{(4)},Y_{\bf 1^{\prime}}^{(8)},Y_{\bf 1}^{(12)}$ with proper $k_{H_{u}}$ and $k_{H_{d}}$). So, we have
$\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1}^{(16)}}}\eta(\omega)^{-6}Y_{\bf 1}^{(16)}(\omega+\epsilon)(\sqrt{3})^{(k_{u}+k_{d})/2}~{},$ (68)
$\approx$ $\frac{\alpha}{360.275}(0.8e^{-0.131i})^{-6}\Lambda_{SUSY}1325.96\epsilon^{4}(\sqrt{3})^{13/2}~{},$
$\approx$ $4.98\times 10^{2}e^{0.785i}\alpha\Lambda_{SUSY}\epsilon^{4}~{}.$
Here $\bar{N}_{Y_{\bf 1}^{(16)}}=360.275$ denotes the global normalization factor Petcov:2023fwh , which does not depend on the VEV of the modulus $\tau$ and defined by
$(\overline{{\rm N}}^{(K)}_{\rm Y})^{2}=\int_{D}\int(N^{(K)}_{\rm Y}(\tau,\tau^{*}))^{2}\,\frac{dxdy}{y^{2}}\equiv\int_{D}\int\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}(2\,y)^{K}\frac{dx\,dy}{y^{2}}\,,$ (69)
with
$N_{Y}^{(K)}=\left(\sum\limits_{i}\left|Y_{i}^{(K)}\right|^{2}(2{\rm Im}(\tau))^{K}\right)^{1/2}~{},$ (70)
or
[…]
for non-cusp modular forms, within which $T$ is a real constant, the region $D_{T}$ is the lower part of the fundamental domain bounded from above by the line $y=T$, $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
For $\epsilon\sim 10^{-5}$ and $\Lambda_{SUSY}$ typically of order the reduced Planck scale, we can obtain the desired $\mu_{eff}$ at electroweak scale. Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be the consequence of SUSY breaking that depends on the choice of $f_{0}$, for example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$. In general, adopting the modular form $Y_{\bf 1}^{(4(k+2))}$ with $(k\in N)$, we can obtain suppression factor $\epsilon^{k+2}$ for tiny $\epsilon(f_{0})$. We show some of the benchmark points that can lead to the suppression factor in Table 5.1.
$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $X_{nl}/M_{\mathrm{Pl}}$ $\epsilon$ 00 $3$ $0.2$ $10.0$ $5.194\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.2$ $8.767\times 10^{-4}$ 00 $2$ $0.2$ $0.1$ $0.02$ $4.227\times 10^{-4}$ 00 $3$ $0.2$ $0.1$ $0.02$ $6.916\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ $0.02$ $9.396\times 10^{-4}$ $1$ $3$ $0.2$ $0.1$ 00 $4.711\times 10^{-4}$ Table 1: Parameter choices for scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-4}|$).
5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
Although the scalar potential for modulus in general diverges for $\tau\rightarrow i\infty$, the introduction of nilpotent Goldstino $X$ superfield or multiple moduli/matter fields may alter the shape of the scalar potential at intermediate values of $\tau$, making it possible the stabilization of modulus at $\tau_{0}$ with medium $\Im\tau_{0}$ (for example, $\Im\tau_{0}\sim 5-10$). Detailed discussions on such a possibility will be given in our subsequent studies. In this work, we just concentrate on the implications of modulus stabilization at medium $\Im\tau$ on the SUSY $\mu$-problem. Note that the stabilized VEV $\tau_{0}$ (with medium $\Im\tau_{0}$) can be regarded to lie close to the $i\infty$ fixed point.
To obtain the suppressed effective $\mu$-parameter, in addition to the suppression factor from powers of $q^{1/24}$ or $(2\Im\tau)^{-1}$, we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$. For example, the modular functions satisfy
$Y_{\bf 1^{\prime}}^{(4)}(i\infty)=Y_{\bf 1^{\prime}}^{(8)}(i\infty)=Y_{\bf 1^{\prime\prime}}^{(8)}(i\infty)=0~{},$ (72)
as well as the vanishing singlet components from their direct products, such as the modular functions $Y_{\bf 1}^{(12)}(i\infty)$,$Y_{\bf 1^{\prime}}^{(16)}(i\infty)$.
The asymptotic behavior of the previous modular forms near $i\infty$ within the fundamental domain can be calculated from Petcov:2023vws (for $z_{0}$ with $\Im z_{0}$ being sufficient large)
$Y_{\bf 1^{\prime}}^{(4)}(z_{0})$ $=$ $-12p\epsilon+96p^{4}\epsilon^{4}~{},$
$Y_{\bf 1^{\prime}}^{(8)}(z_{0})$ $=$ $-6p\epsilon-1230p^{4}\epsilon^{4}~{},$
$Y_{\bf 1^{\prime\prime}}^{(8)}(z_{0})$ $=$ $144p^{2}\epsilon^{2}-2304p^{5}\epsilon^{5}~{},$ (73)
with
$\epsilon=\exp\left[-\frac{2\pi}{3}\Im z_{0}\right]~{},~{}~{}p=\exp\left[i\frac{2\pi}{3}\Re z_{0}\right],$ (74)
where $\epsilon\ll 1$.
Let us suppose that the modulus could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account.
The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (65), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6, is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{u}+k_{d})/2}~{},$ (75)
$\approx$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{r}-k_{3}+3)/2}~{},$
$\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~{}.$
with proper normalization factors for $H_{u},H_{d}$ from the Kahler potential.
For fixed $k_{r}$, we require that the factor
$\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{\frac{k_{3}}{2}}\ll 1.$ (76)
As the factor
$e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}>1~{},$ (77)
for all $t_{0}>0$, the value of $k_{3}$ needs to be negative.
We can choose the following modular weights and representations
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime\prime}},~{}14),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1},~{}15)~{},$ (78)
and the choice of modular forms $Y_{\bf 1^{\prime\prime}}^{(8)}$ (or possibly $Y_{\bf 1}^{(12)}$, $Y_{\bf 1^{\prime}}^{(16)}$ etc with proper $k_{H_{u}}$ and $k_{H_{d}}$) with $k_{3}=-21$ to generate suppressed $\mu_{eff}$ parameter. In this case, the effective $\mu_{eff}$ parameter is
$\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}\Lambda_{SUSY}\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{-21/2}144e^{-\frac{4\pi t_{0}}{3}}(2t_{0})^{11/2}~{},$ (79)
$=$ $4.46{\alpha}\Lambda_{SUSY}e^{-29\pi t_{0}/6}(2t_{0})^{16}~{},$
with the global normalization factor ${\bar{N}_{Y_{\bf 1^{\prime}}^{(8)}}}=32.2794$.
It can be calculated that the value of $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be satisfied for $t_{0}\gtrsim 5.58$. If we choose the modular form from direct product of $(Y_{\bf 1^{\prime}}^{(4)})^{k+2}$ with larger $k\in N$, smaller $t_{0}$ is needed to get electroweak scale $\mu_{eff}$ parameter. Tiny value of $\mu_{eff}$ is an joint effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
6 Conclusions
It is known that the solutions of SUSY $\mu$ problem involve the forbidden of the bare $\mu$-term via some symmetry and the re-generation of the effective $\mu$-term after the symmetry breaking. We propose that, the symmetry that responsible for the solution of the SUSY $\mu$-problem is a modular flavor symmetry. The bare $\mu$-term is forbidden by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations for $H_{u},H_{d}$ superfields. Then, the effective $\mu$-term will be re-generated after the modulus field is stabilized, with the corresponding modulus VEV breaks the modular symmetry completely or partially (at the fixed points).
Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions. The goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into a constrained nilpotent superfield $X_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions, described in the non-linear SUSY realization scheme with nilpotent goldstino $X_{nl}$. The presence of $X_{nl}$ contributions can slightly change the original stabilized modulus VEV to a value located in the vicinity of the fixed points. Natural value of $\mu_{eff}$, which is much suppressed relative to the SUSY scale, can be the consequence of series expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), which is similar to the mechanism to naturally generate mass hierarchies in the vicinity of the fixed points. For modulus VEV with medium $\Im\tau$, the much suppressed $\mu_{eff}$ is the joint consequence of suppression from powers of $q$ [or $(2\Im\tau)^{-1}$], with the exponent of the power determined by the assignments of the modular weights for $H_{u}$ and $H_{d}$, and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
Appendix A Properties of Modular $A_{4}$
The finite modular group $A_{4}\cong\Gamma_{3}$ can generated by
$S^{2}=(ST)^{3}=T^{3}=1~{}.$ (80)
There are 4 inequivalent irreducible representations of $A_{4}$: three singlets $\mathbf{1}$, $\mathbf{1}^{\prime}$, $\mathbf{1}^{\prime\prime}$ and a triplet $\mathbf{3}$.
The triplet representation $\mathbf{3}$ (in the basis where $T$ is diagonal) is given by
$S=\frac{1}{3}\begin{pmatrix}-1~{}&2~{}&2\\ 2~{}&-1~{}&2\\ 2~{}&2~{}&-1\end{pmatrix},~{}\quad~{}T=\begin{pmatrix}1~{}&~{}0~{}&~{}0\\ 0~{}&~{}\omega~{}&~{}0\\ 0~{}&~{}0~{}&~{}\omega^{2}\end{pmatrix}\,.$ (81)
The decompositions of the direct product of $A_{4}$ representations are
$\mathbf{1}^{\prime}\otimes\mathbf{1}^{\prime}=\mathbf{1}^{\prime\prime},~{}~{}~{}\mathbf{1}^{\prime}\otimes\mathbf{1}^{\prime\prime}=\mathbf{1},~{}~{}~{}\mathbf{1}^{\prime\prime}\otimes\mathbf{1}^{\prime\prime}=\mathbf{1}^{\prime}\,,$
$\mathbf{3}\otimes\mathbf{3}=\mathbf{1}\oplus\mathbf{1^{\prime}}\oplus\mathbf{1^{\prime\prime}}\oplus\mathbf{3}_{S}\oplus\mathbf{3}_{A}\,,$ (82)
where $\mathbf{3}_{S(A)}$ denotes the symmetric (antisymmetric) combinations, respectively.
Given two $A_{4}$ triplet representations $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $\beta=(\beta_{1},\beta_{2},\beta_{3})$, the irreducible representations obtained from their direct products are:
$\mathbf{1}=\alpha_{1}\beta_{1}+\alpha_{2}\beta_{3}+\alpha_{3}\beta_{2}\,,$
$\mathbf{1}^{\prime}=\alpha_{3}\beta_{3}+\alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\,,$
$\mathbf{1}^{\prime\prime}=\alpha_{2}\beta_{2}+\alpha_{1}\beta_{3}+\alpha_{3}\beta_{1}\,,$
$\mathbf{3}_{S}=(2\alpha_{1}\beta_{1}-\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2},2\alpha_{3}\beta_{3}-\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1},2\alpha_{2}\beta_{2}-\alpha_{1}\beta_{3}-\alpha_{3}\beta_{1})\,,$
$\mathbf{3}_{A}=(\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2},\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1},\alpha_{3}\beta_{1}-\alpha_{1}\beta_{3})\,.$ (83)
Modular weight $k=2$ and the representation ${\bf r}={\bf 3}$ are given as $Y^{(2)}_{\mathbf{3}}=\left(Y_{1},Y_{2},Y_{3}\right)^{T}$ with
$Y_{1}(\tau)$ $=$ $\frac{i}{2\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}-\frac{27\eta^{\prime}(3\tau)}{\eta(3\tau)}\right],$
$Y_{2}(\tau)$ $=$ $\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega^{2}\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\omega\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right],$
$Y_{3}(\tau)$ $=$ $\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\omega^{2}\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right]\,,$ (84)
where $\eta(\tau)$ is the Dedekind eta-function,
$\eta(\tau)=q^{1/24}\prod_{n=1}^{\infty}(1-q^{n}),~{}~{}~{}q=e^{2\pi i\tau}\,.$ (85)
The $q-$expansions of $Y_{1,2,3}(\tau)$ are given as
$Y_{1}(\tau)=1+12q+36q^{2}+12q^{3}+84q^{4}+72q^{5}+\dots\,,$
$Y_{2}(\tau)=-6q^{1/3}(1+7q+8q^{2}+18q^{3}+14q^{4}+\dots)\,,$
$Y_{3}(\tau)=-18q^{2/3}(1+2q+5q^{2}+4q^{3}+8q^{4}+\dots)\,.$ (86)
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grant numbers Nos.12075213 and 12335005, by the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046.