### v3

# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

Abstract
We propose a solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry, and an effective $\mu$-term is regenerated following the stabilization of the modulus field. We examine the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described by the non-linear SUSY realization scheme involving a nilpotent Goldstino $\textbf{X}_{nl}$ superfield. The inclusion of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near some fixed point. A natural small $\mu_{eff}$, significantly smaller than the SUSY scale, can result from either the expansion of typical modular forms using a small deviation parameter near the fixed point $\omega$, or from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
1 Introduction
Low-energy supersymmetry (SUSY) is well motivated from a theoretical standpoint and is considered to be one of the most compelling extensions of the Standard Model. It not only address the quadratic divergence problem associated with the Higgs mass but also enables successful gauge coupling unification. Notably, the observed 125 GeV scalar, discovered by both the ATLAS ATLAS and CMS CMS collaborations of LHC, falls miraculously within the narrow $115\sim 135$ GeV ’window’ predicted by the minimal supersymmetric standard model (MSSM). This can be regarded as strong supports for low energy SUSY.
Although the framework of low energy SUSY is attractive, it still has several unsatisfactory features. In addition to the lack of conclusive signals for superparticles at the LHC, the MSSM is also bothered by the SUSY $\mu$-problem. The $\mu$-parameter, a SUSY preserving parameter in the superpotential, however, is required (without too much miraculous cancelation) to be roughly of the order of the soft SUSY breaking scale, as determined by the minimization of the MSSM Higgs potential. Many proposals have been put forward to address the $\mu$-problem, for example, by singlet extension of MSSM to next-to-minimal supersymmetric standard model (NMSSM) or by introducing non-renormalizable terms in the superpotential via Kim-Nilles mechanism KN (or in the Kahler potential via Giudice-Masiero mechanism GM ). Most solutions involve two steps Bae:2019dgg : First, the appearance of $\mu$-parameter is prohibited, usually by some discrete or continuous symmetry. Second, the $\mu$ term is regenerated via some mechanism such as symmetry breaking. For detailed realizations of these two steps, one can refer to Bae:2019dgg .
We propose to tackle the $\mu$-problem with a similar but different approach. The symmetry adopted to forbid the $\mu$-parameter is a modular flavor symmetry. The regeneration of $\mu$ term is realized following the modular symmetry breaking, with the modulus stabilization mechanism accompanied by typical SUSY breaking contributions.
Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis. Modular symmetry can naturally arise in string theory, for example, in the (exact) moduli spaces of Calabi-Yau threefolds like the quintic, and more generically in large classes of elliptic Calabi-Yau threefolds. Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the standard model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings. For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt . In the modular flavor symmetry framework, the superpotential is tightly constrained by the modular symmetry, with the relevant Yukawa couplings being modular forms of a certain level $N$, which are functions of a modulus field. The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$). In this framework, the masses and mixings of quarks and leptons can be predicted with very few input parameters. Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; CentellesChulia:2023osj ; Kumar:2023moh ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been used in various studies to explain the low energy flavor structures of the Standard Model. Modular symmetry can also be used to solve the strong CP problem Feruglio:2023uof ; Feruglio:2024ytl . Moreover, modular flavor symmetries can be combined with the GUT framework to further enhance the predictive power of the models, such as combining $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking. It can also be the only source of CP symmetry breaking Novichkov:2019sqv ; Feruglio:2021dte ; Feruglio:2024ytl . The VEV of the modulus field $\tau$ generally break the modular symmetry completely, except at three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain. The modular forms (thus the Yukawa couplings) and the fermion mass matrices are fixed once the modular flavour symmetry is fully or partially broken by the VEV of $\tau$. From a top-down perspective, the value of modulus VEV can be determined dynamically with a modular-invariant scalar potential in 4d $N=1$ supergravity. This is also crucial for giving nonzero masses to the modulus field, rendering it phenomenological acceptable.
Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that are close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy. In the presence of additional ”matter”-like fields, it is possible to obtain a dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induce a SUSY breaking vacuum shift to a tiny positive vacuum energy. The presence of the anti-brane-like sources can be described in terms of a nilpotent Goldstino superfield KS in a supersymmetric effective field theory. In the framework of non-linearly realized SUSY, the couplings of chiral, gauge and gravity superfields to the constrained nilpotent Goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action. Contributions from nilpotent chiral superfield can uplift the scalar potential responsible for the stabilization of the modulus field.
This paper is organized as follows. In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry. In Sec 3, we discuss the solution of SUSY $\mu$-problem using modular flavor symmetry. In Sec 4, we discuss the stabilization of singlet modulus field in the presence of SUSY breaking contributions described by the non-linear realized SUSY scheme with the nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield. In Sec 5, we discuss the regeneration of a naturally small $\mu_{eff}$ after modulus stabilization, when the modulus field is stabilized near some fixed points, such as $\omega\equiv e^{i2\pi/3}$ and $i\infty$ Sec 6 contains our conclusions.
2 Modular $\Gamma_{N}$ symmetry
The full modular group is defined as the set of linear fractional transformations acting on the complex modulus $\tau$ in the upper half complex plane
$\gamma=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right):\tau\longmapsto\gamma\tau=\frac{a\tau+b}{c\tau+d},~{}~{}\text{with}~{}~{}a,b,c,d\in\mathbb{Z},~{}~{}ad-bc=1,~{}~{}~{}{\rm Im}\tau>0\,.$ (1)
This group can be generated by two generators, $S$ and $T$, which act as follows
$S:\tau\mapsto-\frac{1}{\tau},~{}~{}~{}~{}\quad T:\tau\mapsto\tau+1\,,$ (2)
These generators satisfy the relations
$S^{4}=1,~{}~{}(ST)^{3}=1,~{}~{}S^{2}T=TS^{2}.$ (3)
The $S,T$ generators can be represented by the following $SL(2,Z)$ matrices,
$S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~{}~{}~{}\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)\,.$ (4)
Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined as $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, where $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ is given by
$\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}N\geq 3\\ \Gamma(N)/\{\pm 1\},~{}~{}~{}N=2\end{array}\right.~{}~{}~{},$ (7)
with $\Gamma(N)$ the principal congruence subgroup of level $N$
$\Gamma(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL(2,\mathbb{Z}),~{}b=c=0\,(\mathrm{mod~{}N}),a=d=1\,(\mathrm{mod~{}N})\right\}\,.$ (8)
Finite modular groups $\Gamma_{N}$ can be obtained by further imposing the condition $T^{N}=1$ in addition to $S^{2}=(ST)^{3}=1$.
Modular form $f(\tau)$ of weight $2k$ and level $N$ is a holomorphic function of the complex modulus $\tau$ with well-defined transformation properties under the action of ${\Gamma}(N)$
$f\left(\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{2k}f(\tau)~{},~{}~{}~{}\forall~{}\gamma\equiv\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in{\Gamma}(N)\,.$ (9)
Modular forms of weight $2k$ and level $N$ span a linear space of finite dimension. They can be arranged into some modular multiplets $Y_{\mathbf{r}}^{(2k)}$ that transform as irreducible representation $\mathbf{r}$ of the finite modular group $\Gamma_{N}$ up to automorphy factor Feruglio:2017spp , i.e.
$Y_{\mathbf{r}}^{(2k)}(\gamma\tau)=(c\tau+d)^{2k}\rho_{\mathbf{r}}(\gamma)Y_{\mathbf{r}}^{(2k)},~{}~{}~{}~{}\forall~{}\gamma\in\overline{\Gamma}\,,$ (10)
where $\gamma$ is the representative element of the coset $\gamma\overline{\Gamma}(N)$ in $\Gamma_{N}$, and $\rho_{\mathbf{r}}(\gamma)$ is the representation matrix of the element $\gamma$ in the irreducible representation $\mathbf{r}$. The actions of modular group on supermultiplets read
$\phi_{i}(\gamma\tau)=(c\tau+d)^{-k_{\phi}}\rho^{\mathbf{r^{\prime}}}_{ij}(\gamma)\phi_{j}~{},~{}~{}~{}\forall\gamma\in\bar{\Gamma}~{},$ (11)
where $k_{\phi}$ denotes the modular weight of $\phi$, and $\rho$ is a unitary representation of the quotient group $\bar{\Gamma}_{N}$.
3 Natural $\mu$-term in modular flavor model
As noted in Bae:2019dgg , to solve the SUSY $\mu$ problem, the appearance of a $\mu$ term in the superpotential should be prohibited by some symmetry. We propose that the superfields $H_{u}$, $H_{d}$ transform non-trivially under modular transformations, thereby prohibiting the ordinary $\mu H_{u}H_{d}$ term by modular symmetry.
The relevant Lagrangian can be given as
${\cal L}$ $\supseteq$ $\int d^{4}\theta\left[-3\log(2\Im\tau)+\frac{H_{u}^{\dagger}H_{u}}{(2\Im\tau)^{k_{u}}}+\frac{H_{d}^{\dagger}H_{d}}{(2\Im\tau)^{k_{d}}}+\cdots\right]$ (12)
$+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
where the coefficient for bilinear $H_{u}H_{d}$ term is given by
$\tilde{\mu}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ to ensure the modular invariance of
$e^{G}\equiv e^{K}|W|^{2}~{}.$ (14)
Here, $\alpha\sim{\cal O}(1)$ is a dimensionless parameter and $\Lambda_{SUSY}$ is typically of order the SUSY scale. The Dedekind $\eta$-function is a modular form with modular weight $1/2$ under modular transformation up to a phase. Its transformation under $\gamma$ is given by
$\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~{},~{}~{}\forall\gamma\in\overline{\Gamma}~{},$ (15)
where $\epsilon_{\eta}(\gamma)$ depends on $\gamma$ (but not on $\tau$) and can be expressed in terms of the Rademacher $\Phi(\gamma)$ function
$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right],$ (16)
that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx .
The expansion of the $\eta(\tau)$ function is
$\eta(\tau)$ $=$ $q^{1/24}\sum\limits_{n=-\infty}^{\infty}(-1)^{n}q^{n(3n-1)/2},$ (17)
$=$ $q^{1/24}\left[1+\sum\limits_{n=1}^{\infty}(-1)^{n}\left(q^{n(3n-1)/2}+q^{n(3n+1)/2}\right)\right],$
$=$ $q^{1/24}\left[1-q-q^{2}+q^{5}+q^{7}-q^{12}-\cdots\right],~{}~{}~{}~{}~{}q\equiv e^{i2\pi\tau}~{}.$
It is evident that the Kahler potential for superfields with non-trivial modular weights, such as $H_{u}$ and $H_{d}$ superfields, is non-minimal. Therefore, after the modulus field $\tau$ acquires a VEV through the modulus stabilization mechanism, the $H_{u},H_{d}$ fields need to be normalized properly to get their canonical kinetic forms. This normalization corresponds to the following replacements
$H_{u}\rightarrow H_{u}(2\Im\tau)^{k_{H_{u}}/2}~{},~{}~{}~{}H_{d}\rightarrow H_{d}(2\Im\tau)^{k_{H_{d}}/2}~{}.$ (18)
Thus, when the modulus field is stabilized, the effective $\mu$-parameter is regenerated as
$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{H_{u}}+k_{H_{d}})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
We need to explain why $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and of order the electroweak scale or soft SUSY breaking scale $m_{soft}$. We will show that, the smallness of $\mu_{eff}$ is the consequence of large suppression factors, either originate from some power of deviation parameter that characterizes the deviation of modulus VEV from typical fixed point, or from non-trivial power (determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of a small factor related to the modulus VEV. All these suppression factors are determined by the modulus stabilization mechanism, with the stabilized modulus VEVs being affected by SUSY breaking effects.
4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
To stabilize the modulus field, we can utilize effective supergravity scalar potentials derived from a weight $-3$ superpotential to determine the minimum. The modulus VEV can always be restricted to the fundamental domain
${D}_{H}\equiv\left\{\tau\in\mathcal{H}:-\frac{1}{2}\leq\Re{\tau}<\frac{1}{2},\,|\tau|>1\right\}\cup\left\{\tau\in\mathcal{H}:-\frac{1}{2}<\Re\tau\leq 0,\,|\tau|=1\right\}\,,$ (20)
of the modular group, as any choice of the modulus VEV in the upper-half plane can be uniquely related to a $\tau\in D_{H}$ in the fundamental domain via a modular transformation, which is physically equivalent to such a choice. Discussions on modulus stabilization with single modulus and multiple moduli can be found in Kobayashi:2019xvz ; Kobayashi:2020hoc ; Kobayashi:2020uaj ; Ishiguro:2020tmo ; NPP:2201.02020 ; PLNSR:2304.14437 ; King:2023snq ; King:2024ssx ; Ding:2024neh ; Higaki:2024pql ; Higaki:2024jdk ; Kobayashi:2023spx ; Funakoshi:2024yxg .
To ensure the modular invariance and Kahler transformation invariance of
$e^{G}\equiv e^{K}|W|^{2}~{},$ (21)
the superpotential for the modulus should take the form
$W(\tau)=c_{1}^{3}\frac{H(\tau)}{\eta(\tau)^{6}},$ (22)
with the Kahler potential
$K_{mod}=-3\log(2\Im\tau)~{},$ (23)
in the unit $M_{Pl}=1$. Here, $c_{1}M_{Pl}$ denotes the typical mass scale for the superpotential, where $c_{1}$ is a dimensionless number.
The modular invariant holomorphic function $H(\tau)$ is given by
$H(\tau)_{m,n}$ $=$ $\left(j(\tau)-1728\right)^{m/2}j(\tau)^{n/3}{\cal P}\left(j(\tau)\right),$ (24)
$\equiv$ $\left(\frac{G_{6}(\tau)}{\eta(\tau)^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\eta(\tau)^{12}}\right)^{n}{\cal P}\left(j(\tau)\right)~{},$
where ${\cal P}\left(j(\tau)\right)$ denotes a polynomial with respect to $j(\tau)$. The Klein $j$-function is defined as a modular form of zero weight
$j(\tau)$$\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\eta(\tau)^{24}},$ (25)
$\approx$ $744+q^{-1}+196884q+21493760q^{2}+864299970q^{3}+{\cal O}(q^{4})~{},$
with $j(i\infty)$ diverges, $j(\omega)=0$ and $j(i)=1728$.
We should note that, in such a stabilization mechanism, the fixed points $z=i,\omega\equiv e^{i2\pi/3}$ are always the extrema of the scalar potential, although they may not be the global minimum. With proper SUSY breaking contributions, possibly from the ”matter”-like sector, the stabilized modulus value, which located at/near the fixed points $z_{FP}=\omega,i$, can slightly deviate from the original global minimum. The minimum of the scalar potential can also possibly be uplifted to dS (de Sitter) type. In fact, the classical 10d supergravity Maldacena-Nunez no-go result MN suggests that any constructions leading to a dS vacuum must include non-classical ingredients, such as D-branes, O-planes or quantum corrections. To describe the scenario in which SUSY is linearly realized in a gravitational sector and non-linearly realized in the uplifting SUSY breaking sector (for example, from anti-D branes), the formalism of non-linearly realized SUSY as formulated by KS can be used for such a system. It was proved in Burgess:2021juk that, with the non-linear SUSY realization scheme, the supergravity form is stable even if the supermultiplets in SUSY breaking sector are split badly enough to allow some of its members to be integrated out while the others are not.
When SUSY is spontaneously broken in supergravity, the Goldstino is eaten by the gravitino via the super-Higgs mechanism. When such a breaking happens at low energies compared with the Planck scale, the Goldstino couplings can be conveniently described by a (constrained) independent superfield in the supergravity effective action. The Goldstino provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. The emergence of the constraints may also be viewed as an effective field theory limit in which certain mass parameters become very large and the corresponding modes decouple.
The coupling of non-linearly realized SUSY with a nilpotent chiral Goldstino superfield, subjected to constraints ${\textbf{X}}_{nl}^{2}=0$, can be described by
${\cal L}\supseteq\int d^{4}\theta\textbf{X}_{nl}^{\dagger}\textbf{X}_{nl}+\left[\int d^{2}\theta f_{0}\textbf{X}_{nl}+h.c.\right]~{}.$ (26)
The nilpotent constraints for the chiral Goldstino superfield can be solved to be
$\textbf{X}_{nl}=\frac{\psi_{X}\psi_{X}}{2F_{X}}+\sqrt{2}\theta\psi_{X}+\theta^{2}F_{X}~{},$ (27)
where all the fields are the functions of $y^{\mu}\equiv x^{\mu}+i\theta\sigma^{\mu}\bar{\theta}$. The lowest component of $\textbf{X}_{nl}$ is eliminated by the nilpotent constraints $\textbf{X}_{nl}^{2}=0$ and is no longer an independent complex scalar field. This leaves only the Goldstino field $\psi_{X}$ and an independent auxiliary field $F_{X}$ in $\textbf{X}_{nl}$ superfield. The Goldstino field $\psi_{X}$, which is the fermionic component of $\textbf{X}_{nl}$, eventually mixes with the gravitino to give it a mass through the super-Higgs mechanism. The lowest component of $\textbf{X}_{nl}$, which is denoted as ${X}_{nl}$ for the Goldstino bilinear ${\psi_{X}\psi_{X}}$, obviously satisfies ${X}_{nl}^{2}=0$ due to its anti-commuting Grassman variable nature. The chiral superfield $\textbf{X}_{nl}$ can be seen as the long-distance expression of the microscopic superfield X that satisfies $\bar{D}^{\dot{\alpha}}{\cal J}_{\alpha\dot{\alpha}}=D_{\alpha}\textbf{X}$ for Ferrara-Zumino multiplet ${\cal J}_{\alpha\dot{\alpha}}$ KS , which can be used even when SUSY is spontaneously broken and can be followed consistently along the renormalization group flow down to the deep IR. As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $\textbf{X}_{nl}$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent Goldstino superfield. From the perspective of low energy effective field theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for discussions on modulus stabilization and vacuum uplifting.
When coupled to supergravity, the most general form (at the two-derivative level) for the action in superspace is given by
[…]
with
$K=\textbf{X}_{nl}^{\dagger}\textbf{X}_{nl}~{},~{}~{}~{}W=f_{0}\textbf{X}_{nl}+c_{1}^{3}~{},$ (29)
where without loss of generality we take $f_{0}$ and $c_{1}$ to be real. Integrating out the auxillary field $F_{X}$ gives $F_{X}=-(W_{\textbf{X}_{nl}})^{*}=-f_{0}$. To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
$K$ $\supseteq$ $-3\log(2\Im\tau)+\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}~{},$
$W$ $\supseteq$ $\left(f_{0}\textbf{X}_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
The supergravity scalar potential can be derived from
$V=e^{K}\left[K^{\alpha\bar{\beta}}\nabla_{\alpha}W\nabla_{\bar{\beta}}W-3W\overline{W}\right],$ (31)
where the Kahler-covariant derivative of the superpotential is
$\nabla_{\alpha}W=W_{\alpha}+K_{\alpha}W~{}.$ (32)
From the Kahler potential, we can calculate the inverse Kahler metric
$K^{{\alpha}\bar{\beta}}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}\textbf{X}_{nl}}{3}\\ -\frac{i(2\Im\tau)k_{X}\textbf{X}^{\dagger}_{nl}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}(\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl})}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
Here the indices $\alpha,\beta\sim(\tau,\textbf{X}_{nl})$ and the nilpotency of $\textbf{X}_{nl}$ is used in the calculation.
The scalar potential for the modulus field $\tau$ and $X_{nl}$ can be calculated as
$V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (36)
$=$ $\frac{e^{\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left(1-\frac{k_{X}}{3}\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}\right)\left|\tilde{W}_{M}\tilde{c}_{M}+H(\tau)\tilde{W}_{T}\right|^{2}\right.$
$+$ $(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}|X_{nl}|^{2}}{3(2\Im\tau)^{k_{X}-1}}\right]\left|H(\tau)\tilde{W}_{X}\right|^{2}$
$+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}(\tilde{W}_{X}H(\tau))^{*}\left(\tilde{W}_{M}\tilde{c}_{M}+H(\tau)\tilde{W}_{T}\right)\right]$
-- $\left.3\left|\tilde{c}_{W}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
where
$\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
$\tilde{W}_{T}$$\equiv$ $k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\left[\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}+\frac{ic_{1}^{3}X_{nl}^{*}}{f_{0}(2\Im\tau)^{k_{X}+1}\eta(\tau)^{2k_{X}}}\right]~{},$
$\tilde{c}_{M}$ $=$ $f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}~{},$
$\tilde{W}_{X}$ $=$ $f_{0}\eta(\tau)^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{2(\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (37)
The constraint $X_{nl}^{2}=0$ for the lowest component of $\textbf{X}_{nl}$ is used to eliminate some of the high order $X_{nl}$ terms.
We also use the following expression for the derivative of Dedekind $\eta$ function
$\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}~{},$ (38)
where the holomorphic Eisenstein series $G_{2}$, which is a quasi-modular form
$G_{2}(\gamma\tau)=(c\tau+d)^{2}G_{2}(\tau)-2\pi ic(c\tau+d)~{},$ (39)
is related to its weight two non-holomorphic counterpart $\hat{G}_{2}(\tau,\bar{\tau})$ as
$\hat{G}_{2}(\tau,{\tau}^{*})=G_{2}(\tau)-\frac{\pi}{\Im\tau}~{}.$ (40)
The generic complicate form of the scalar potential may obscure the discussions on the stabilization mechanism. Therefore, we would like to concentrate on its typical simplification forms
  • Scenario with $k_{X}=0$:
    When $k_{X}=0$, the scalar potential simplifies significantly.
    The potential can then be written as
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (41)
    $=$ $\frac{e^{|X_{nl}|^{2}}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\tilde{c}_{M}\right|^{2}\right.$
    $+$ $\left.\left|H(\tau)\right|^{2}\left|f_{0}\left(1+{|X_{nl}|^{2}}\right)+{{X}^{*}_{nl}c_{1}^{3}}\right|^{2}-3\left|\tilde{c}_{M}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right.$
    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}{|X_{nl}|^{2}}+2\Re\left[f_{0}c_{1}^{3}{X}_{nl}^{*}\right]+|X_{nl}|^{2}c_{1}^{6}\frac{}{}\right)~{}$
    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}|X_{nl}|^{2}+c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\right\}~{}$
    $+$ $\frac{|X_{nl}|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{},$
    where
    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
    $\tilde{c}_{M}$ $=$ $f_{0}X_{nl}+c_{1}^{3}~{}.$ (42)
    Note the constraints $X_{nl}^{2}=({X}^{*}_{nl})^{2}=0$ are used in the scalar potential.
    In this scenario, the modulus VEV can be determined by the extreme conditions of the scalar potential
    $\frac{\partial}{\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=\frac{\partial}{\partial\tau^{*}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})=0~{},$
    and the minimum condition can be determined by the positive definiteness of Hessian matrix in terms of
    $\frac{\partial^{2}}{\partial x^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}+\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
    $\frac{\partial^{2}}{\partial y^{2}}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $2\left(\frac{\partial^{2}}{\partial\tau\partial\tau^{*}}-\Re\frac{\partial^{2}}{\partial\tau\partial\tau}\right)V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$
    $\frac{\partial^{2}}{\partial x\partial y}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ $=$ $-2\Im\frac{\partial^{2}}{\partial\tau\partial\tau}V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*}),$ (43)
    where $\tau\equiv x+iy\equiv\rho_{\tau}e^{i\theta_{\tau}}$. It was argued in target:space ; Ding:2024neh ; Higaki:2024pql that the modular invariance and $\tau\leftrightarrow-\tau^{*}$ symmetry of the scalar potential ensure that the first derivative of scalar potential along certain directions at the boundary of the fundamental domain vanishes
    $\left.\frac{\partial}{\partial x}V\right|_{x=\pm 1/2,0}=0~{},~{}~{}~{}\left.\frac{\partial}{\partial\rho_{\tau}}V\right|_{\rho_{\tau}=1}=0~{}.~{}~{}~{}$ (44)
    We then determine the extreme conditions
    $\frac{\partial}{\partial\rho}V=\frac{\partial}{\partial\theta}V=0~{},$ (45)
    for $X_{nl}\equiv\rho e^{i\theta}$. We have the following discussions for some non-generic simplified scenarios:
    • Case $\rho=0$:
      This corresponds to $X_{nl}=0$, reducing to the case with only a single $\tau$ modulus field, which has already been extensively discussed target:space ; NPP:2201.02020 .
    • Case $f_{0}=0$:
      The extrema of the scalar potential are located at the parameter points with $X_{nl}=0$. Again, the modulus VEV of the corresponding extrema is the same as the case with only a single $\tau$ modulus field.
    • Case $c_{1}=0$:
      The extrema of the scalar potential are also located at the parameter point with $X_{nl}=0$, and the corresponding potential energy is
      $\left.V\right|_{X_{nl}=0}=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left|{H(\tau)}\right|^{2}f_{0}^{2}\geq 0~{},$ (46)
      for values of $\tau$ within the fundamental region. To obtain a vanishing (very tiny positive) vacuum energy $\Lambda^{4}\sim 10^{-120}M_{Pl}^{4}\approx 0$, we can find the parameter points that are solutions of ${H(\tau)}=0$ and also satisfy the minimum condition (with respect to $\tau$ and $\tau^{*}$) with $\tilde{W}_{M}=0$. It can be checked that solutions of ${H(\tau)}=0$ can also satisfy the minimum condition $\tilde{W}_{M}=0$, which will be clear shortly. Therefore, the stabilized modulus VEV is located at the solutions of ${H(\tau)}=0$ with (almost) vanishing vacuum energy in this case.
    We now focus on the general case with $f_{0},c_{1},X_{nl}\neq 0$. It can be calculated that the extreme condition for $\theta$ is
    $\frac{\partial}{\partial\theta}V=-\frac{2f_{0}c_{1}^{3}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\rho\sin\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]=0.$ (47)
    First, we should note that, as $[(2\Im\tau)|\eta(\tau)|^{4}]^{-1}$ diverges when $\Im\tau=\infty$, the $\tau=i\infty$ point is not a solution of the extreme condition for $\theta$.
    For generic $\theta$ (with $\theta\neq 0,\pi$ to render $X_{nl}$ not real) in the case $f_{0},c_{1},\rho\neq 0$, the extreme condition for $\theta$ reduces to the constraint
    $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{}.$ (48)
    After substituting this constraint in the extreme condition for $\rho$, we have
    $\frac{\partial}{\partial\rho}V=4\rho f_{0}^{2}\frac{\left|H(\tau)\right|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}=0~{}.$ (49)
    Such an extreme condition for $\rho$ can be satisfied with
    $\left|H(\tau)\right|^{2}=0~{}.$ (50)
    Thus, we have the extreme conditions for the case with $f_{0},c_{1}\neq 0$ and a non-vanishing non-real $X_{nl}$
    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{}.$ (51)
    Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes. Therefore, the presence of a non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe in this case.
    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires
    $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0,~{}~{}{\rm or}~{}~{}~{}~{}{\cal P}(j(\tau))=0.$ (52)
    Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively. As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, any complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all unless we specify the concrete form of ${\cal P}(j(\tau))$.
    Then, we would like to focus on the constraint
    $\tilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0~{},$ (53)
    which satisfies $H(\tau)=0$. Such a constraint reduces to
    $H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))\right\}^{\prime}=0~{},$ (54)
    $\Rightarrow$ $j^{\prime}(\tau)\left[\frac{n}{3}j(\tau)^{n/3-1}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))+\frac{m}{2}j(\tau)^{n/3}[j(\tau)-1728]^{m/2-1}{\cal P}(j(\tau))\right.$
    $\left.+j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{}.$
    So $H^{\prime}(\tau)=0$ can trivially hold when $H(\tau)=0$.
    On the other hand, $H^{\prime}(\tau)=0$ can also be satisfied by the other two conditions
    $\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$ (55)
    or $j^{\prime}(\tau)=0$.
    We have the following discussions for the condition (55)
    • $m+n\neq 0$:
      When $n\neq 0$ and $m=0$, condition (55) reduces to
      $\frac{d{\cal P}(j(\tau))}{{\cal P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)}~{},$ (56)
      which gives
      ${\cal P}(j(\tau))=C_{0}\left[j(\tau)\right]^{-n/3}~{}.$ (57)
      Unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$. Similarly, when $m\neq 0$ and $n=0$, we can deduce that $m\in 2\mathbb{Z}$ should be negative. For $m\neq 0$ and $n\neq 0$, similar deduction can give
      ${\cal P}(j(\tau))=C_{0}^{\prime}\left[j(\tau)\right]^{-n/3}\left[{j(\tau)-1728}\right]^{-m/2}~{},$ (58)
      which also requires both $m\in 2\mathbb{Z}$ and $n\in 3\mathbb{Z}$ being negative to guarantee ${\cal P}(j(\tau))$ being a polynomial of $j(\tau)$. As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
    • $m=0$ and $n=0$:
      With the constraints $m=n=0$, the condition (55) reduces to
      $\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}=0~{},$ (59)
      which gives
      ${\cal P}(j(\tau))={\rm Constant}.$ (60)
      As $H(\tau)=0$, this constant should be zero.
    Therefore, the condition (55) requires a vanishing $H(\tau)$. For the condition $j^{\prime}(\tau)=0$, from the derivative of the Klein $j$-function
    $\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~{},$ (61)
    we can see that its solutions are $\omega$ and $i$, leading to no new additional constraints. Therefore, the extreme conditions (51) can reduce to the requirement of vanishing $H(\tau)$.
    Before we conclude this scenario, it should be noted that $\theta=0,\pi$, which corresponds to real $X_{nl}=\pm\rho$ with positive $\rho$, are also solutions of the the extreme condition (47) for $\theta$. In this case, the scalar potential reduces to
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (62)
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right.$
    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}\rho^{2}\pm 2f_{0}c_{1}^{3}\rho+\rho^{2}c_{1}^{6}\frac{}{}\right)~{}$
    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right\}~{}$
    $+$ $\frac{\rho^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{}.$
    The scalar potential, which is a quadratic polynomial of $\rho$, can be written as
    $V=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left[C_{1}\rho^{2}\pm C_{2}\rho+C_{3}\right]~{},$ (63)
    where
    $C_{1}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}(f_{0}^{2}+c_{1}^{6})-2\left|H(\tau)\right|^{2}c_{1}^{6}~{},$
    $C_{2}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}2f_{0}c_{1}^{3}-4\left|H(\tau)\right|^{2}f_{0}c_{1}^{3}~{},$
    $C_{3}$ $=$ $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}\left(f_{0}^{2}-3c_{1}^{6}\right)~{}.$ (64)
    The scalar potential has a minimum with the corresponding $\rho$ parameter
    $\rho_{0}=\mp\frac{C_{2}}{2C_{1}}~{},$ (65)
    and the potential energy
    $\left.V\right|_{\rho=\rho_{0}}=C_{3}-\frac{C_{2}^{2}}{2C_{1}}~{}.$ (66)
    The modulus VEV is determined by the minimum conditions for the potential (66) with respect to $\tau,\tau^{*}$, and the condition $C_{2}^{2}=2C_{1}C_{3}$ should be established to obtain a vanishing (almost vanishing tiny positive) vacuum energy. It can be checked from previous discussions that solutions of ${H(\tau)}=0$ can also satisfy such minimum conditions.
    Thus, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential are located at the points determined by $H(\tau)=0$, and the corresponding vacuum is of Minkowski type. As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are always the local extrema unless $m=n=0$. Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$. Therefore, other minima, which are determined by (44) to lie at the boundary of the fundamental domain or at the imaginary axis, can also be possible.
    In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices. It can be seen that the potential energies are non-negative when $f_{0}\neq 0$.
    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$. We adopt $\mathcal{P}(j)=1$ for the left panel and $\mathcal{P}(j)=j-4.5764069879884376\times 10^{-11}$ for the right panel.
  • Scenario with small $X_{nl}$ (when $|X_{nl}|\rightarrow 0$):
    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the constraint $X_{nl}^{2}=0$. Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below.
    The scalar potential can be simplified to
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (67)
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\left[\eta(\tau)\right]^{2k_{X}}\right]\right)\right.$
    $+$ $\left.\frac{(2\Im\tau)^{2}}{3}\left(2\Re\left[k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\tilde{W}_{E}\right]\right)H(\tau)c_{1}^{3}\tilde{W}_{M}^{*}\right]$
    $+$ $(2\Im\tau)^{k_{X}}\left|H(\tau)\right|^{2}\left(f_{0}^{2}\left|\eta(\tau)\right|^{4k_{X}}+2\Re\left[f_{0}\eta(\tau)^{2k_{X}}\frac{{X}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}\right]\right)$
    $+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}\left(f_{0}[\eta(\tau)]^{2k_{X}}H(\tau)\right)^{*}\tilde{W}_{M}c_{1}^{3}\right]$
    -- $\left.3\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\frac{}{}\right\}~{},$
    with
    $\tilde{W}_{M}$$\equiv$ $H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})~{},$
    $\tilde{W}_{E}$$\equiv$ $\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}~{}.$ (68)
    It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$. Therefore, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point, can possibly be the minimum. The initially stabilized modulus VEV without contributions from $X_{nl}$ sector can then be destabilized by new contributions from non-vanishing $X_{nl}$. We will discuss the extrema conditions for the most general case in our subsequent studies.
It should be noted that the modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous, which is independent on the following discussions on the regeneration of small $\mu_{eff}$ term following modulus stabilization.
5 Regeneration of naturally small $\mu_{eff}$ following modulus stabilization
[…]
It is known that the fixed points $\tau_{FP}=i,\omega,i\infty$ for modular theories based on the $\Gamma_{N}$ invariance can preserve $Z_{2}^{S},Z_{3}^{ST},Z_{N}^{T}$ residual symmetries, respectively Novichkov:2021evw . We can infer from the modular invariance of scalar potential that the fixed points should be the local extrema of the scalar potential, as they act non-trivially (with residual symmetries) on the derivatives of the potential and must vanish. Although they may not be the global minima, experiences from modulus stabilization in string theory indicate that the minima should always lie at/near $\tau_{FP}=i,\omega$ for the tree-level scalar potential, as the scalar potential generally diverges for $\tau\rightarrow i\infty$.
It was found in Novichkov:2021evw ; Feruglio:2021dte that the hierarchies among the charged leptons and quarks mass can follow solely from the properties of the modular forms (strictly speaking, the representation of the mass term bilinear with respect to the residual group) in the vicinity of symmetric points, avoiding the fine-tuning of the constant parameters (or without the need to introducing extra weighton-like scalar fields). Similar ideas, augmented with additional suppression powers of $q^{1/24}$ (or $(2\Im\tau)^{-1}$), can be used to generate the hierarchy between $\Lambda_{SUSY}$ and $\mu_{eff}$.
Studies on target-space modular symmetry in string theory indicates target:space that all extrema lie on the boundary of the fundamental region or on the imaginary axis. We anticipate that, due to SUSY breaking contributions (for example, from $X_{nl}$ sector), the extrema can change with respect to the change of $f_{0}$ so as that the VEV $\tau_{0}$ can be altered slightly away from the fixed points $\tau_{FP}=i,\omega$. Such deviations from fixed points by SUSY breaking effects can lead to large suppression factors that regenerate $\mu_{eff}$ of order the electroweak scale.
Besides, although the scalar potential for modulus field $\tau$ in general diverges when $\tau\rightarrow i\infty$, the presence of $X_{nl}$ contributions can alter the scalar potential, making it possible to stabilize $\tau$ at an intermediate value. Consequently, for an appropriate choice of $k_{3}$, the effective $\mu_{eff}$ parameter can naturally be much smaller than the SUSY scale $\Lambda_{SUSY}$, due to a large suppression factor originating from both powers of $q$ [or $(2\Im\tau)^{-1}$] and typical suppression behavior of modular forms away from the fixed point $i\infty$ (at $\tau_{0}$ with medium $\Im\tau_{0}$).
To regenerate the effective $\mu$-term with a stabilized modulus field, we need to construct a modular weight $-3$ bilinear for $H_{u}$ and $H_{d}$ using proper modular forms. For a modular invariant bilinear of the form $\psi_{i}^{c}M_{ij}(\tau)\psi_{j}$, where $M(\tau)_{ij}$ is a modular form of level $N$ and weight $K$, the zero entries of $M(\tau)_{ij}$ at a symmetric point in the mass matrices will generically become non-zero when $\tau$ deviates slightly from the symmetric point. The magnitudes of these residual symmetry breaking entries are controlled by the size of the departure $\epsilon$ from $\tau_{sym}$ and the field transformation properties under the residual symmetry group, which may depend on the modular weights Petcov:2023vws ; Novichkov:2021evw ; Petcov:2022fjf . That is, the entries of the mass matrices in the vicinity of the symmetric point are subjected to corrections of ${\cal O}(\epsilon^{l})$, where the powers $l$ corresponding to the representations of the residual symmetry group and depends only on how the representations of the fermion fields in the mass term bilinear decompose under the considered residual symmetry group. The degree of suppression, given by the integer $l$, can take values $l=0,1,\cdots,N-1$ in the case of $\tau_{sym}=i\infty$, $l=0,1,2$ in the case of $\tau_{sym}=\omega$ and $l=0,1$ in the case of $\tau_{sym}=i$. In Novichkov:2021evw , the authors systematically classify the residual symmetry representations for different choices of $\Gamma_{N}$ representations of matter fields.
Such a formalism can also be applied to the regeneration of small $\mu_{eff}$ for the $H_{u}H_{d}$ bilinear to achieve a desired suppression factor ${\cal O}(\epsilon^{l})$ of the deviation parameter $\epsilon$ (from a given symmetric point). In our case, we need to specify the representations of the Higgs superfields $H_{u},H_{d}$ in MSSM under the residual symmetry group $\mathbb{Z}^{T}_{N}$ (with $N=3$) for $\tau_{sym}=i\infty$ and the residual group $\mathbb{Z}^{ST}_{3}$ for $\tau_{sym}=\omega$. Obviously, $H_{u},H_{d}$ should always be singlets under the modular flavor symmetry $\Gamma_{N}$. They can only be assigned some singlet representations of $\Gamma_{N}$ with non-trivial modular weights. The decompositions of the representation $({\bf r};k)$ multiplet of $\Gamma_{N}^{\prime}$ under the residual symmetry can be found in the appendix of Novichkov:2021evw . So, the desired suppression factor for the regenerated $\mu_{eff}$ can be directly obtained with proper choice of singlet representations and modular weights.
5.1 Natural $\mu_{eff}$ with modulus stabilization near the fixed point $\omega\equiv e^{i2\pi/3}$
In the vicinity of $\omega$, the degree of suppression for a bilinear is proven Novichkov:2021evw to be given by ${\cal O}(\epsilon^{l})$ with $l$ taking values $l=0,1,2$. In this study, we choose to directly calculate the suppression factor and its coefficient for the bilinear $H_{u}H_{d}$ using explicit model construction involving appropriate modular forms. It should be emphasized again that the exponentiation power of deviation parameter in the suppression factor depends only on the $\mathbb{Z}^{ST}_{3}$ (or $\mathbb{Z}^{T}_{3}$ for modular VEVs near $i\infty$) charge of $H_{u},H_{d}$ instead of the explicit form of modular forms.
For modular $A_{4}$ symmetry, at the fixed point $\omega$, the following modular forms satisfy
$Y_{\bf 1}^{(4)}(\omega)=Y_{\bf 1}^{(8)}(\omega)=Y_{\bf 1^{\prime}}^{(8)}(\omega)=0~{}.$ (70)
The small deviation of modulus VEV away from $\omega$, induced by the modulus stabilization mechanism, can act as the desired suppression parameter for $\mu_{eff}$.
Expressions of modular forms with larger modular weights can be constructed from $Y_{\bf 3}^{(2)}\equiv(Y_{1},Y_{2},Y_{3})$, for example,
$Y_{\bf 1^{\prime}}^{(8)}$ $=$ $(Y_{1}^{2}+2Y_{2}Y_{3})(Y_{3}^{2}+2Y_{1}Y_{2})~{},~{}~{}~{}Y_{\bf 1}^{(8)}=(Y_{1}^{2}+2Y_{2}Y_{3})^{2}~{}.$ (71)
From the expansion Petcov:2022fjf of modular functions at the point $z_{0}=\omega+\epsilon$ (with $\epsilon$ depending on the choices of $f_{0}$ such that $|\epsilon|\ll 0$)
$\frac{Y_{2}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $\omega\left(1+2.235i\epsilon-2.95717\epsilon^{2}+\cdots\right)~{},$
$\frac{Y_{3}(z_{0})}{Y_{1}(z_{0})}$ $\simeq$ $-\frac{1}{2}\omega^{2}\left(1+4.47i\epsilon-10.9096\epsilon^{2}+\cdots\right),$ (72)
we can obtain the expansion for $Y_{\bf 1}^{(8)}(z_{0})$ etc as
$Y_{\bf 1}^{(8)}(\omega+\epsilon)$ $\simeq$ $-36.4138\epsilon^{2}-259.129i\epsilon^{3}+\cdots~{},$ (73)
with $Y_{1}(\omega)\approx 0.948674$.
[…]
$W\supseteq\alpha\Lambda_{SUSY}\frac{Y_{\bf r}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}~{},$ (74)
where $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ and ${\bf r}(H_{u})\otimes{\bf r}(H_{d})\otimes{\bf r}\supset{\bf 1}$. When the modulus field is stabilized close to the point $\omega$, the effective $\mu$-term is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega+\epsilon)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(2\Im\tau)^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (75)
$=$ $\alpha\Lambda_{SUSY}\eta^{-2k_{3}}(\omega)Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{(k_{r}-k_{3}+3)/2}$
$\approx$ $\alpha\Lambda_{SUSY}\left[0.711506\cdot e^{-0.523599i}\right]^{-k_{3}/2}Y_{\bf r}^{(k_{r})}(\omega+\epsilon)(\sqrt{3})^{k_{r}+3/2}~{}.$
To achieve a significantly suppressed $\mu_{eff}$, we either need a very large negative value $k_{3}$ or a tiny $\epsilon$. However, calculations show the suppression from $k_{3}$ power is not very efficient for obtaining a tiny number. Therefore, we choose to use a small $\epsilon$ to suppress $\mu_{eff}$ with appropriate modular weights and representations for $H_{u},H_{d}$.
For instance, we can choose
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime}},~{}4),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime\prime}},~{}4)~{},$ (76)
and the modular forms $Y_{\bf 1}^{(8)}$ (or, for example, $Y_{\bf 1}^{(4)}$ with suitable $k_{H_{u}}$ and $k_{H_{d}}$). Thus, we have
$\mu_{eff}$ $=$ $\frac{\alpha\Lambda_{SUSY}}{|\bar{N}_{Y_{\bf 1}^{(8)}}|}\eta(\omega)^{-6}Y_{\bf 1}^{(8)}(\omega+\epsilon)(\sqrt{3})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (77)
$\approx$ $-0.927\times 10^{3}e^{i\frac{\pi}{4}}\alpha\Lambda_{SUSY}\epsilon^{2}~{}.$
Here $|\bar{N}_{Y_{\bf 1}^{(8)}}|=1.3423$ denotes the global normalization factor Petcov:2023fwh for $Y_{\bf 1}^{(8)}$, which is independent of the modulus VEV and is defined by
$(\overline{{\rm N}}^{(K)}_{\rm Y})^{2}=\int_{D}\int(N^{(K)}_{\rm Y}(\tau,\tau^{*}))^{2}\,\frac{dxdy}{y^{2}}\equiv\int_{D}\int\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}(2\,y)^{K}\frac{dx\,dy}{y^{2}}\,,$ (78)
with
$N_{Y}^{(K)}=\left(\sum\limits_{i}\left|Y_{i}^{(K)}\right|^{2}(2{\rm Im}(\tau))^{K}\right)^{1/2}~{},$ (79)
or
[…]
for non-cusp modular forms, where $T$ is a real constant, $D_{T}$ is the region below the line $y=T$ in the fundamental domain with $T>1$, and $|a_{0}|^{2}$ is the constant $q$-independent term in the $q$-expansion of $\sum_{i}|Y^{(K)}_{{\bf r}i}(\tau)|^{2}$.
From the decompositions of $\Gamma_{N}$ representations under the residual symmetry group in appendix of Novichkov:2021evw , it can be verified that the suppression factor of $\epsilon$ for the $H_{u}(~{}{\bf 1^{\prime}},~{}4)$ and $H_{d}(~{}{\bf 1^{\prime\prime}},~{}4)$ bilinear is given by $({k_{H_{u}}+k_{H_{d}}+3})~{}{\rm modulo}~{}3=2$, which is consistent with the result from our explicit construction.
For $\epsilon\sim 10^{-5}$, $\alpha\sim{\cal O}(1)$ and $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV for $m_{3/2}\sim{\cal O}({\rm 10^{2}~{}GeV})$ in gravity mediation SUSY breaking mechanism, we can obtain a desired $\mu_{eff}$ at electroweak scale. If $\Lambda_{SUSY}$ is naively taken to be the Planck scale, smaller $\epsilon\sim 10^{-10}$ or a tiny $\alpha$ is needed to achieve an electroweak-scale $\mu_{eff}$. Note that the smallness of deviation parameter $\epsilon$ (away from $\omega$) can be a consequence of SUSY breaking, which depends on the choice of $f_{0}$. For example, if the parameter choice with $f_{0}=0$ leads to the global minimum at $\omega$, SUSY breaking contributions may slightly shift the minimum to some point near $\omega$. We present some benchmark points that can lead to small suppression factors in Table 5.1.
$m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $\epsilon$ $\mathcal{P}(j(\tau))$ 00 $3$ $0.2$ $0.01$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.529\times 10^{-5}$ $1$ 00 $2$ $0.2$ $0.1$ $6.758\times 10^{-6}$ $1$ 00 $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ $1$ $3$ $0.2$ $0.1$ $4.711\times 10^{-5}$ $1$ 00 $3$ $0.2$ $0.1$ $2.920\times 10^{-5}$ $j-4146.6830+184.0955i$ $1$ $3$ $0.2$ $0.1$ $6.587\times 10^{-5}$ $j-4144.8675+185.3946i$ Table 1: Parameter choices for the scenario with $k_{X}=0$, which can stabilize the modulus at the point $\omega+\epsilon$ (for $|\epsilon\sim 10^{-5}|$).
5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
When the modulus field is stabilized near $i\infty$, similar to the case where modulus VEV lies in the vicinity of $\omega$, the exponentiation power $l$ of deviation parameter $\epsilon$ in the suppression factor for the $H_{u}H_{d}$ bilinear can also be proven Novichkov:2021evw to depend only on the $\mathbb{Z}^{T}_{3}$ charge of $H_{u},H_{d}$, rather than the explicit form of modular forms. Since the coefficient of $\epsilon^{l}$ cannot be simply determined and may involve additional suppression factors, we opt to directly calculate the suppression factor and the coefficient for the bilinear $H_{u}H_{d}$ using an explicit model construction.
[…]
To obtain the suppressed effective $\mu$-parameter, in addition to possible suppression factor from powers of $q^{1/24}$ (from $\eta(\tau)$) or $(2\Im\tau)^{-1}$ (from normalization factor), we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$. For example, the modular functions satisfy
$Y_{\bf 1^{\prime}}^{(4)}(i\infty)=Y_{\bf 1^{\prime}}^{(8)}(i\infty)=Y_{\bf 1^{\prime\prime}}^{(8)}(i\infty)=0~{},$ (81)
as well as the vanishing singlet components from their direct products, such as the modular function $Y_{\bf 1}^{(12)}(i\infty)$.
The asymptotic behavior of the previous modular forms near $i\infty$ within the fundamental domain can be calculated from Petcov:2023vws (for $z_{0}$ with $\Im z_{0}$ sufficiently large)
$Y_{\bf 1^{\prime}}^{(4)}(z_{0})$ $=$ $-12p\epsilon+96p^{4}\epsilon^{4}+{\cal O}(\epsilon^{7})~{},$
$Y_{\bf 1^{\prime}}^{(8)}(z_{0})$ $=$ $-6p\epsilon-1230p^{4}\epsilon^{4}+{\cal O}(\epsilon^{7})~{},$
$Y_{\bf 1^{\prime\prime}}^{(8)}(z_{0})$ $=$ $144p^{2}\epsilon^{2}-2304p^{5}\epsilon^{5}+{\cal O}(\epsilon^{8})~{},$ (82)
with
$\epsilon=\exp\left[-\frac{2\pi}{3}\Im z_{0}\right]~{},~{}~{}p=\exp\left[i\frac{2\pi}{3}\Re z_{0}\right],$ (83)
where $\epsilon\ll 1$.
Let us suppose that the modulus field could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account. Naively, if the desired suppression factor $10^{-20}$ to regenerate $\mu_{eff}$ comes solely from $\epsilon^{2}$ when $\Lambda_{SUSY}$ is naively taken to be the Planck scale, very large $t_{0}$ (here $t_{0}\gtrsim 10$) is needed. Such a large $t_{0}$ possibly causes difficulties for modulus stabilization realizations, as the scalar potential for modulus field generally diverges for $\tau\rightarrow i\infty$. Therefore, it is desirable to introduce additional suppression factors other than $\epsilon$ so that medium large $t_{0}$ can lead to desired suppression factors.
The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (74), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$, is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (84)
$\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~{},$
where proper normalization factors for $H_{u}$,$H_{d}$ from the Kahler potential are taken into account.
For fixed $k_{r}$, we require that the factor
$\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{\frac{k_{3}}{2}}\ll 1,$ (85)
is tiny to act as an additional suppression factor other than $\epsilon$. As the factor
$e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}>1~{},$ (86)
for all $t_{0}>0$, the value of $k_{3}$ needs to be negative.
We can choose the following modular weights and representations
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1},~{}16),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime}},~{}16)~{},$ (87)
and the choice of modular form $Y_{\bf 1^{\prime\prime}}^{(8)}$ (or, for example, $Y_{\bf 1}^{(12)}$, $Y_{\bf 1^{\prime}}^{(16)}$ etc with proper $k_{H_{u}}$ and $k_{H_{d}}$) with $k_{3}=-21$ to generate a suppressed $\mu_{eff}$ parameter. In this case, the effective $\mu_{eff}$ parameter is
$\mu_{eff}$ $=$ $\frac{\alpha}{\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}\Lambda_{SUSY}\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{-21/2}144e^{-\frac{4\pi t_{0}}{3}}(2t_{0})^{11/2}~{},$ (88)
$=$ $7.05{\alpha}\Lambda_{SUSY}e^{-29\pi t_{0}/6}(2t_{0})^{16}~{},$
where ${\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}=20.4152$ is the global normalization factor for the modular form $Y_{\bf 1^{\prime\prime}}^{(8)}$.
Calculations show that the value $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be achieved for $t_{0}\gtrsim 5.58$, if $\Lambda_{SUSY}$ is naively taken to be the Planck scale. For $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV, it is sufficient to choose $t_{0}\gtrsim 3.2$ to satisfy $\mu_{eff}\lesssim 10^{-8}M_{SUSY}$. The tiny value of $\mu_{eff}$ results from the combined effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
It can also be verified that the suppression power of $\epsilon=\exp[-\frac{2\pi}{3}\Im z_{0}]$ for $H_{u}(~{}{\bf 1},~{}16)$ and $H_{d}(~{}{\bf 1^{\prime}},~{}16)$ bilinear is given by the power factor $l$ of $(\rho_{i}^{c}\rho_{j})^{*}=\zeta^{l}$ for $\zeta=\exp(2\pi i/3)$, which can be read out from Novichkov:2021evw to be $(1\cdot\zeta)^{*}=\zeta^{2}$ and predict $l=2$. This power factor for $\epsilon$ is consistent with the result from our explicit construction.
6 Conclusions
In this study, we propose a novel solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations to the $H_{u},H_{d}$ superfields. The effective $\mu$-term is then regenerated following the stabilization of the modulus field, with the corresponding modulus VEV breaks the modular symmetry partially (at the fixed points) or completely (in the vicinity of the fixed points). Since naturalness requires $\mu_{eff}$ to lie at the electroweak scale or soft SUSY breaking scale, a large suppression factor leading to $\mu_{eff}\ll\Lambda_{SUSY}$ is required when the modulus field is stabilized. Such a large suppression factor can naturally arise from a modulus VEV in the vicinity of certain fixed points.
Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions. The Goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into the constrained nilpotent superfield $\textbf{X}_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. We discuss the stabilization mechanism of a single modulus field in the presence of SUSY breaking contributions, described by the non-linear SUSY realization scheme with a nilpotent Goldstino $\textbf{X}_{nl}$. The presence of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near the fixed points.
Natural value of $\mu_{eff}$, which is significantly suppressed relative to the SUSY scale, can be achieved from the expansion of typical modular forms using a tiny deviation parameter near the fixed points. This is similar to the mechanism that can naturally generate mass hierarchies for quarks and leptons in the vicinity of certain fixed points. The exponentiation power of the deviation parameter in the suppression factor, in fact, depends solely on the $\mathbb{Z}_{3}^{ST}$ charge (for fixed point $\omega$), or $\mathbb{Z}^{T}_{3}$ charge (for fixed point $i\infty$) of $H_{u}$, $H_{d}$, rather than the explicit form of modular forms.
For a modulus VEV with an intermediate value of $\Im\tau$, the significantly suppressed $\mu_{eff}$ results from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$], where the exponentiation power determined by the modular weights of $H_{u}$ and $H_{d}$, along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
The modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous. The regeneration of small $\mu_{eff}$ term following modulus stabilization, which just needs a modulus VEV near some fixed points, is independent on the concrete modulus stabilization mechanism.
Appendix A Properties of Modular $A_{4}$
The finite modular group $A_{4}\cong\Gamma_{3}$ can generated by
$S^{2}=(ST)^{3}=T^{3}=1~{}.$ (89)
There are 4 inequivalent irreducible representations of $A_{4}$: three singlets $\mathbf{1}$, $\mathbf{1}^{\prime}$, $\mathbf{1}^{\prime\prime}$ and a triplet $\mathbf{3}$.
The triplet representation $\mathbf{3}$ (in the basis where $T$ is diagonal) is given by
$S=\frac{1}{3}\begin{pmatrix}-1~{}&2~{}&2\\ 2~{}&-1~{}&2\\ 2~{}&2~{}&-1\end{pmatrix},~{}\quad~{}T=\begin{pmatrix}1~{}&~{}0~{}&~{}0\\ 0~{}&~{}\omega~{}&~{}0\\ 0~{}&~{}0~{}&~{}\omega^{2}\end{pmatrix}\,.$ (90)
The decompositions of the direct product of $A_{4}$ representations are
$\mathbf{1}^{\prime}\otimes\mathbf{1}^{\prime}=\mathbf{1}^{\prime\prime},~{}~{}~{}\mathbf{1}^{\prime}\otimes\mathbf{1}^{\prime\prime}=\mathbf{1},~{}~{}~{}\mathbf{1}^{\prime\prime}\otimes\mathbf{1}^{\prime\prime}=\mathbf{1}^{\prime}\,,$
$\mathbf{3}\otimes\mathbf{3}=\mathbf{1}\oplus\mathbf{1^{\prime}}\oplus\mathbf{1^{\prime\prime}}\oplus\mathbf{3}_{S}\oplus\mathbf{3}_{A}\,,$ (91)
where $\mathbf{3}_{S(A)}$ denotes the symmetric (antisymmetric) combinations, respectively.
Given two $A_{4}$ triplet representations $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $\beta=(\beta_{1},\beta_{2},\beta_{3})$, the irreducible representations obtained from their direct products are:
$\mathbf{1}=\alpha_{1}\beta_{1}+\alpha_{2}\beta_{3}+\alpha_{3}\beta_{2}\,,$
$\mathbf{1}^{\prime}=\alpha_{3}\beta_{3}+\alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\,,$
$\mathbf{1}^{\prime\prime}=\alpha_{2}\beta_{2}+\alpha_{1}\beta_{3}+\alpha_{3}\beta_{1}\,,$
$\mathbf{3}_{S}=(2\alpha_{1}\beta_{1}-\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2},2\alpha_{3}\beta_{3}-\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1},2\alpha_{2}\beta_{2}-\alpha_{1}\beta_{3}-\alpha_{3}\beta_{1})\,,$
$\mathbf{3}_{A}=(\alpha_{2}\beta_{3}-\alpha_{3}\beta_{2},\alpha_{1}\beta_{2}-\alpha_{2}\beta_{1},\alpha_{3}\beta_{1}-\alpha_{1}\beta_{3})\,.$ (92)
Modular weight $k=2$ and the representation ${\bf r}={\bf 3}$ are given as $Y^{(2)}_{\mathbf{3}}=\left(Y_{1},Y_{2},Y_{3}\right)^{T}$ with
$Y_{1}(\tau)$ $=$ $\frac{i}{2\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}-\frac{27\eta^{\prime}(3\tau)}{\eta(3\tau)}\right],$
$Y_{2}(\tau)$ $=$ $\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega^{2}\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\omega\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right],$
$Y_{3}(\tau)$ $=$ $\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega\frac{\eta^{\prime}((\tau+1)/3)}{\eta((\tau+1)/3)}+\omega^{2}\frac{\eta^{\prime}((\tau+2)/3)}{\eta((\tau+2)/3)}\right]\,,$ (93)
where $\eta(\tau)$ is the Dedekind eta-function,
$\eta(\tau)=q^{1/24}\prod_{n=1}^{\infty}(1-q^{n}),~{}~{}~{}q=e^{2\pi i\tau}\,.$ (94)
The $q-$expansions of $Y_{1,2,3}(\tau)$ are given as
$Y_{1}(\tau)=1+12q+36q^{2}+12q^{3}+84q^{4}+72q^{5}+\dots\,,$
$Y_{2}(\tau)=-6q^{1/3}(1+7q+8q^{2}+18q^{3}+14q^{4}+\dots)\,,$
$Y_{3}(\tau)=-18q^{2/3}(1+2q+5q^{2}+4q^{3}+8q^{4}+\dots)\,.$ (95)
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grant numbers Nos.12075213 and 12335005, by the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046.